# Identifying Weight-Variant Latent Causal Models

### v6

@@ A.7 Detailed Results on fMRI
Figure 21 shows the recovered latent six signals (Blue) and the true ones (Red) within one day by the proposed SuaVE.
Figure 21: Recovered latent six signals (Blue) and the true ones (Red) within one day by the proposed SuaVE.
The recovered latent fMRI variables within one day by iVAE are depicted by Figure 22.
Figure 22: Recovered latent six signals (Blue) and the true ones (Red) within one day by iVAE.
The recovered latent fMRI variables within one day by VAE are depicted by Figure 23.
Figure 23: Recovered latent six signals (Blue) and the true ones (Red) within one day by VAE.
The recovered latent fMRI variables within one day by $\beta$-VAE are depicted by Figure 24.
The recovered latent fMRI variables within one day by CausalVAE are depicted by Figure 25.
Figure 24: Recovered latent six signals (Blue) and the true ones (Red) within one day by $\beta$-VAE.
Figure 25: Recovered latent six signals (Blue) and the true ones (Red) within one day by CausalVAE.
Implementation Details: For fMRI data, again, we used the same network architecture for encoder (e.g. 3-layer fully connected network with 30 hidden nodes for each layer) and decoder (e.g. 3-layer fully connected network with 30 hidden nodes for each layer) parts in all these models. For prior model in the proposed SuaVE and iVAE, we use 3-layer fully connected network with 30 hidden nodes for each layer. We assign an 3-layer fully connected network with 30 nodes to generate the weights to model the relations among latent causal variables in the proposed SuaVE. For hyper-parameters, we set $\beta=4,25,50$ for $\beta$-VAE. For CausalVAE, we use the hyper-parameters setting as recommended by the authors (Yang et al., 2021), for generating observed $\mathbf{x}$, we use an invertible 2-layer multi-layer perceptron on the fMRI with different 5 random seeds.

### v7

@@ A.7 Detailed Results on fMRI
Figure 21: Recovered six signals (Blue) and the true ones (Red) within one day by SuaVE.
Figure 22: Recovered six signals (Blue) and the true ones (Red) within one day by iVAE.
Figure 23: Recovered six signals (Blue) and the true ones (Red) within one day by VAE.
Figure 24: Recovered six signals (Blue) and the true ones (Red) within one day by $\beta$-VAE.
Figure 25: Recovered six signals (Blue) and the true ones (Red) by CausalVAE.
Figures 21-25 show the recovered latent six signals (Blue) and the true ones (Red) within one day by the proposed SuaVE, iVAE, VAE, $\beta$-VAE, and CausalVAE, respectively. Regarding to implementation details, for fMRI data, again, we used the same network architecture for encoder (i.e., 3-layer fully connected network with 30 hidden nodes for each layer) and decoder (i.e., 3-layer fully connected network with 30 hidden nodes for each layer) parts in all these models. For prior model in the proposed SuaVE and iVAE, we use 3-layer fully connected network with 30 hidden nodes for each layer. We assign an 3-layer fully connected network with 30 nodes to generate the weights to model the relations among latent causal variables in the proposed SuaVE. For hyper-parameters, we set $\beta=4$, 25, 50 for $\beta$-VAE. For CausalVAE, we use the hyper-parameters setting as recommended in Yang et al. (2021), for generating observed $\mathbf{x}$, we use an invertible 2-layer multi-layer perceptron on the fMRI with different 5 random seeds.